find the equation of the lines that are parallel and perpendicular to the line 7x+2y=-4 and pass through the point (6,-3)
Answer by mananth(16949)

7x+2y=-4
2y=-7x-4
y=-7x/2 -2
slope = -7/2
PARALLEL LINE
slope will be the same = -7/2 point ( 6,-3)
find b
y = mx +b
-3 = -7/2 * 6 +b
-3=-21+b
b=18
the equation will be
y = -7x/2 + 18
...
PERPENDICULAR
slope = 2/7 point = (6,-3)
Find b
y=mx + b
-3 = 2/7 * 6 +b
-3-12/7 =b
(-21-12)/7 =b
-33/7 = b
y = 2x/7-33/7
...
m.ananth@hotmail.ca